A piece of wire 8 feet in length is to be cut into two pieces. Each piece will then be bent into a square. Where should the cut be made if the sum of the areas of the squares is to be 2 square feet?
Answer by checkley71(8403)

lets cut the wire in half 4 ft & 4 ft.
bending each 4 ft. length gives us a set of squares with sides=1 ft each
area of each square=1*1=1 ft^2.
since we have 2 of these we have 2*1=2 ft^2.